BP1: Determine normal and transverse shear stress on a simply supported rectangular 5-ft beam with a cross-sectional height oLoad: q=100lb/in
Answer #1

=11 g=loodblin y=111 AK X=4 o - 511 TFB d=112 K L=5ft - Shear force at ct (where, fc = FA-2X = 21 - 2x FA=FB = Fc = 12/1ooxs